if a15 = 20, a30 = -13 then what is r, S50 (Geometric Sequence)
Answer by Theo(13342)

the formula for An of a geometric sequence is:

An = A1 * r ^ (n-1)

the formula for Sn of a geometric sequence is:

Sn = A1 * (1 - r ^ n) / (1 - r)

we first need to find A1 and r and then we can find Sn.

you are given that:

A15 = 20
A30 = -13

if you know that An = A1 * r ^ (n-1), and you replace A1 with Am, then it follows that the formula should then read:

An = Am * r ^ (n-m)

therefore, you get:

A30 = A15 * r ^ (30 - 15)

this becomes A30 = A15 * r ^ (15)

since A30 = -13 and A15 = 20, you get -13 = 20 * r ^ (15)

divide both sides of this equation by 20 to get -13 /20 = r ^ (15)

take the 15th root of both sides of this equation to get (-13 / 20) ^ (1/15) = r

solve for r to get r = -.9716896058

now that you know r, you can solve for A1.

the formula of An = A1 * r ^ (n-1) becomes A15 = A1 * r ^ 14 when n = 15.

since A15 = 20 and r = -.9716896058, this formula then becomes 20 = A1 * -.9716896058 ^ 14.

divide both sides of this equation by -.9716896058 ^ 14 and you get 20 /(-.9716896058 ^ 14) = A1.

solve for A1 to get A1 = 29.89814172

now that you know A1 and r, you can solve for S50.

the formula for Sn is Sn = A1 * (1 - r^n) / (1 - r)

when n = 50 and A1 = 29.89814172 and r = -.9716896058, that formula becomes S50 = 29.89814172 * (1 - (-.9716896058) ^ 50) / (1 - (-.9716896058).

solve for S50 to get S50 = 11.55640593

i took the trouble to do the actual n by n calculations for you in excel so you could see the progression.

this is what it looks like.

$$$